Search Previous Year Questions

Find real questions from previous competitive exam papers with answers and solutions.

Leave the keyword blank to browse saved questions for your selected exam, newest papers first. Add a keyword, year or topic to narrow the results.

Browse complete papers
Showing 10 question(s)
Clear

Question 1

NIMCET 2020• Full Test

If \(\rm \begin{pmatrix} 15\ 8\end{pmatrix}+\begin{pmatrix} 15\ 7\end{pmatrix}=\begin{pmatrix} \rm n\ \rm r\end{pmatrix}\)

then the values of n and r are:

Where, \(\begin{pmatrix} \rm n\ \rm r\end{pmatrix} = {}^nC_r\)

  1. AOption (a)

    16 and 7.

  2. BOption (b)

    16 and 8.

  3. COption (c)

    16 and 9.

  4. DOption (d)

    30 and 15.

View answer & explanation

Correct answer

Answer: (b)B.

16 and 8.

Why this answer is correct

Concept:

  • nCr = nCn-r.
  • If nCx = nCy, then x = y or x + y = n.
  • nCr + nCr-1 = n+1Cr.

 

Calculations:

\(\rm \begin{pmatrix} 15\ 8\end{pmatrix}+\begin{pmatrix} 15\ 7\end{pmatrix}=\begin{pmatrix} \rm n\ \rm r\end{pmatrix}\)

⇒ 15C8 + 15C7 = nCr

Using nCr + nCr-1 = n+1Cr:

⇒ ​15C8 + 15C8-1 = 15+1C8 = nCr

⇒ ​16C8 = nCr

⇒ n = 16 and r = 8.

Question 2

NIMCET 2020• Full Test

In a class of 50 students, it was found that 30 students read "Hitavad", 35 students read "Hindustan" and 10 read neither. How many students read both "Hitavad" and "Hindustan" newpapers?

  1. AOption (a)

    25

  2. BOption (b)

    35

  3. COption (c)

    15

  4. DOption (d)

    30

View answer & explanation

Correct answer

Answer: (a)A.

25

Why this answer is correct

Concept:

Let A and B denote two sets of elements.

  • n(A) and n(B) are the number of elements present in set A and B respectively.
  • n(A ⋃ B) is the total number of elements present in either set A or B.
  • n(A ⋂ B) is the number of elements present in both the sets A and B.
  • n(A ⋃ B) = n(A) + n(B) - n(A ⋂ B)

 

Calculations:

Let A be the set of students who read "Hitavad" and B the set of students who read "Hindustan".

From the given information:

n(A ⋃ B) = 50 - 10 = 40.

n(A) = 30.

n(B) = 35.

By using n(A ⋃ B) = n(A) + n(B) - n(A ⋂ B), we get:

40 = 30 + 35 - n(A ⋂ B)

⇒ n(A ⋂ B) = 25.

∴ The number of students who read both "Hitavad" and "Hindustan" newpapers is n(A ⋂ B) = 25.

Question 3

NIMCET 2020• Full Test

If x = {4n - 3n - 1: n ϵ N} and Y = {9(n - 1) : n ϵ N}, where N is the set of natural numbers, then

  1. AOption (a)

    X ⊂ Y

  2. BOption (b)

    X \(\subseteq\) Y

  3. COption (c)

    X \(\supset\) Y

  4. DOption (d)

    X \(\supseteq\) Y

View answer & explanation

Correct answer

Answer: (a)A.

X ⊂ Y

Why this answer is correct

Concept:

We know that a set of which all the elements are contained in another set is known as a subset

Calculations:

Given, X = {4n - 3n - 1: n ϵ N}, where N is the set of natural numbers so N =1, 2, 3, ....

After putting all these values of N, we will get the elements of X.

⇒ X = {0, 9, 54,....}

Similarly, Y = {9(n - 1) : n ϵ N}, where N is the set of natural numbers,

⇒ Y = {0, 9, 18, 27,....}

Here, all elements of X is contained in set Y. 

Hence, X is a subset of Y.

\(\rm X \subset Y\)

Hence, if x = {4n - 3n - 1: n ϵ N} and Y{9(n - 1) : n ϵ N}, where N is the set of natural numbers, then \(\rm X \subset Y\).

Question 4

NIMCET 2020• Full Test

If A = {x, y, z}, then the number of subsets in powerset of A is

  1. AOption (a)

    6

  2. BOption (b)

    8

  3. COption (c)

    7

  4. DOption (d)

    9

View answer & explanation

Correct answer

Answer: (b)B.

8

Why this answer is correct

Concept:

  • The power set (or powerset) of a Set A is defined as the set of all subsets of the Set A including the Set itself and the null or empty set.

Calculations:

Given, A = {x, y, z}. 

The power set (or powerset) of a Set A is defined as the set of all subsets of the Set A including the Set itself and the null or empty set.

Powerset of A = {ϕ,{x}, {y}, {z}, {x, y}, {y, z}, {x, z},{x, y, z}}.

Hence, the number of subsets in the powerset of A is 8.

Question 5

NIMCET 2020• Full Test

How many words starting with letter D can be formed by taking all letters from word DELHI, so that the letters are not repeated?

  1. AOption (a)

    4

  2. BOption (b)

    12

  3. COption (c)

    24

  4. DOption (d)

    120

View answer & explanation

Correct answer

Answer: (c)C.

24

Why this answer is correct

Concept:

Basic Principle of Counting:

If there are m ways for happening of an event A, and corresponding to each possibility there are n ways for happening of event B, then the total number of different possibilities for happening of events A and B are:

  • Either event A OR event B alone = m + n.
  • Both event A AND event B together = m × n.
  • A number is divisible by 2 if its units digit is either 0, 2, 4, 6 or 8.

 

Permutations/Arrangements:

  • The number of ways in which r objects can be arranged in n places is nPr = \(\rm \dfrac{n!}{(n - r)!}\).
  • The number of ways in which n objects can be arranged in n places is nPn = n!.
  • n! = 1 × 2 × 3 × ... × n.
  • 0! = 1.

 

Calculations:

The number of letters in the word DELHI = 5.

Since, D has to be the first letter, the number of ways in which the first letter can be written is 1.

The remaining 4 letters can be written in any order/arrangement of the remaining 4 letters (ELHI).

∴ The number of ways of writing the 4 letters = 4! = 4 × 3 × 2 × 1 = 24.

∴ The total number of ways of writing the 5 letters = 1 × 24 = 24.

Question 6

NIMCET 2020• Full Test

Naresh has 10 friends, and he wants to invite 6 of them to a party. How many times will 3 particular friends never attend the party?

  1. AOption (a)

    8

  2. BOption (b)

    7

  3. COption (c)

    720

  4. DOption (d)

    35

View answer & explanation

Correct answer

Answer: (b)B.

7

Why this answer is correct

Concept:

\(\rm ^nC_r = \dfrac{n!}{r!(n-r)!}\)

Calculations:

Given, Naresh has 10 friends, and he wants to invite 6 of them to a party.

Remove the 3 particular friends and invite 6 friends from the remaining = 10 - 3 = 7 friends.

This can be done in 7C6 ways

 

Hence, required number of ways = 7C6 

⇒required number of ways = \(\rm \dfrac{7!}{(7-6)!6!}\)

⇒required number of ways = 7 

Hence, Naresh has 10 friends, and he wants to invite 6 of them to a party. total number of ways that 3 particular friends never attend the party = 7

Question 7

NIMCET 2020• Full Test

There is a young boy's birthday party which 3 friends have attended. The mother has arranged 10 games where a prize is awarded for winning a game. The prizes are identical. If each of the 4 children receives at least one prize, then how many distributions of prizes are possible?

  1. AOption (a)

    80

  2. BOption (b)

    84

  3. COption (c)

    70

  4. DOption (d)

    72

View answer & explanation

Correct answer

Answer: (b)B.

84

Why this answer is correct

Concept:

  • The total number of ways in which n objects can be distributed in r people if any person can get 'any number of objects' = n+r-1Cr-1.
  • The total number of ways in which n objects can be distributed in r people if every person must get 'at least one object' = n-1Cr-1.
  • nCr = \(\rm \dfrac{n!}{r!(n-r)!}\).
  • n! = 1 × 2 × 3 × ... × n.
  • 0! = 1.

 

Calculations:

Here, the number of objects is n = 10 and the number of people is r = 4.

Since, every person must get at least one object, the total number of possible distributions is:

n-1Cr-1 = 10-1C4-1 = 9C3 

= \(\rm \dfrac{9!}{3!(9-3)!}=\dfrac{7\times8\times9}{3\times2\times1}\) 

= 84.

Question 8

NIMCET 2020• Full Test

Three cities A, B and C are equidistant from each other. A motorist travels from A to B at 30 km/hr, from B to C at 40 km/hr and from C to A at 50 km/hr. Then the average speed is:

  1. AOption (a)

    39 km/hr

  2. BOption (b)

    40 km/hr

  3. COption (c)

    38.3 km/hr

  4. DOption (d)

    37.6 km/hr

View answer & explanation

Correct answer

Answer: (c)C.

38.3 km/hr

Why this answer is correct

Concept:

  • If the 'Speed' is constant, then the 'Distance' and the 'Time' are directly proportional to each other. It means that, if one of the quantities becomes n times, the other will also become n times.
  • Speed = \(\rm \dfrac{Distance}{Time}\).
  • Average Speed = \(\rm \dfrac{Total\ Distance}{Total\ Time}\).

 

Calculations:

Let's say that the distance between each pair of cities is 600 km.

From A to B:

Distance = 600 km.

Time = \(\rm \dfrac{Distance}{Speed}=\dfrac{600}{30}\) = 20 hrs.

From B to C:

Distance = 600 km.

Time = \(\rm \dfrac{Distance}{Speed}=\dfrac{600}{40}\) = 15 hrs.

From C to A:

Distance = 600 km.

Time = \(\rm \dfrac{Distance}{Speed}=\dfrac{600}{50}\) = 12 hrs.

For overall journey:

Total distance = 600 + 600 + 600 = 1800 km.

Total time = 20 + 15 + 12 = 47 hrs.

Average Speed = \(\rm \dfrac{Total\ Distance}{Total\ Time}=\dfrac{1800}{47}\) = 38.29 km/hr.

Alternate Method 

Formula used:

Speed = \(\rm \dfrac{Distance}{Time}\).

Average Speed = \(\rm \dfrac{Total\ Distance}{Total\ Time}\).

Calculation:

Let's say that the distance between each pair of cities be x km

According to the question,

Distance = x km.

Time = \(\rm \dfrac{Distance}{Speed}=\dfrac{x}{30}\) 

Time = \(\rm \dfrac{Distance}{Speed}=\dfrac{x}{40}\) 

Time = \(\rm \dfrac{Distance}{Speed}=\dfrac{x}{50}\)

Total distance = x/30 + x/40 + x/50 = 47x/600

Average Speed = Total distance/Total time = 3x × 600/47x = 38.29 km/hr.

Question 9

NIMCET 2020• Full Test

A problem in Mathematics is given to 3 students A, B and C. If the probability of A solving the problem is \(\dfrac{1}{2}\) and B not solving it is \(\dfrac{1}{4}\) and the whole probability of the problem being solved is \(\dfrac{63}{64}\), then what is the probability of solving it by C?

  1. AOption (a)

    \(\dfrac18\)

  2. BOption (b)

    \(\dfrac1{64}\)

  3. COption (c)

    \(\dfrac78\)

  4. DOption (d)

    \(\dfrac12\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\dfrac78\)

Why this answer is correct

Concept:

  • P(not E) = 1 - P(E).
  • Probability of a Compound Event [(A and B) or (B and C)] is calculated as:

P[(A and B) or (B and C)] = [P(A) × P(B)] + [P(C) × P(D)]

('and' means '×' and 'or' means '+')

 

Calculations:

It is given that P(A) = \(\dfrac{1}{2}\) and P(not B) = \(\dfrac{1}{4}\). Let's say that the probability of C solving the problem is P(C) = x.

∴ Probability of A not solving the problem = P(not A) = 1 - P(A) = \(1-\dfrac{1}{2}\) = \(\dfrac{1}{2}\).

And, probability of C not solving the problem = P(not C) = 1 - P(C) = 1 - x.

Now, the probability that the problem is not solved at all:

= P(not A) AND P(not B) AND P(not C)

= P(not A) × P(not B) × P(not C)

= \(\rm \dfrac{1}{2}\times\dfrac{1}{4}\times(1-x)\)

= \(\rm \dfrac{1-x}{8}\).

And, the probability that the problem is solved = 1 - Probability that the problem is not solved at all = \(\rm \dfrac{63}{64}\).

⇒ \(\rm 1-\dfrac{1-x}{8}=\rm \dfrac{63}{64}\)

⇒ \(\rm \dfrac{1-x}{8}=1-\rm \dfrac{63}{64}=\dfrac{1}{64}\)

⇒ \(\rm 1-x=\dfrac{1}{8}\)

⇒ \(\rm x=1-\dfrac{1}{8}=\dfrac78\).

∴ The probability of C solving the problem is P(C) = x = \(\dfrac78\).

Question 10

NIMCET 2020• Full Test

A and B play a game where each is asked to select a number from 1 to 25. If the two numbers match, both win a prize. The probability that they will not win a prize in a single trial is

  1. AOption (a)

    \(\frac{1}{25}\)

  2. BOption (b)

    \(\frac{24}{25}\)

  3. COption (c)

    \(\frac{2}{25}\)

  4. DOption (d)

    \(\frac{3}{25}\)

View answer & explanation

Correct answer

Answer: (b)B.

\(\frac{24}{25}\)

Why this answer is correct

Concept:

The probability that they will not win a prize in a single trial = 1 - The probability that they will win a prize in a single trial

Calculations:

Given, A and B play a game where each is asked to select a number from 1 to 25. If the two numbers match, both win a prize.

Total number of possibilities = \(\rm 25 \times 25\)

Favourable number of their wining = 25

⇒ Probability of winning the prize = \(\dfrac {1}{25}\)

The probability that they will not win a prize in a single trial = 1 - The probability that they will win a prize in a single trial

The probability that they will not win a prize in a single trial = 1 - \(\dfrac {1}{25}\)

⇒The probability that they will not win a prize in a single trial = \(\dfrac {24}{25}\)

Previous
Page 1